OpenStudy (anonymous):

A circle has its center at (-2, -3) and a radius of 3 units. What is the equation of the circle?

OpenStudy (anonymous):

(x + 2)2 + (y + 3)2= 9 (x + 2)2 + (y - 3)2= 9 (x - 2)2 + (y - 3)2 = 3 (x + 2)2 + (y + 3)2 = 3

OpenStudy (anonymous):

general equation of circle with center at (a,b) and radiu s r: \[(x-a)^{2}+(y-b)^{2}=r ^{2}\] can you finish alone?

OpenStudy (anonymous):

?????

OpenStudy (anonymous):

(x + 2)2 + (y + 3)2= 9

OpenStudy (anonymous):

lol thanksss